40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 17 16:50:53 2022, Paco, Update, BHD, XARM AUX alignment 
    Reply  Fri Mar 18 17:15:19 2022, Paco, Update, BHD, XARM AUX alignment cont. Screen_Shot_2022-03-18_at_4.09.38_PM.png
       Reply  Mon Mar 21 18:00:05 2022, Ian MacMillan, Update, BHD, XARM AUX alignment cont. PXL_20220322_015247138.jpg
Message ID: 16732     Entry time: Thu Mar 17 16:50:53 2022     Reply to this: 16735
Author: Paco 
Type: Update 
Category: BHD 
Subject: XARM AUX alignment 

[Ian, Paco]

We opened the ITMX chamber and

  • Inspected PR2 and LO1 and didn't see anything suspicious.
  • Took the EQ stops off ITMX in preparation for alignment


  • Ian started at the ETMX station, slightly tweaking the M1 M2 XAUX mirrors by hand and Paco on the ITMX chamber looking at the beam. The beam was visibly making it to the ITMX optic and reflecting back though at a negative pitch angle. No small motion on the M1-M2 knobs caused visible correction on this.
  • We decided to try adding an alignment offset to ITMX, and quickly saturated to 25000 (on the SUS screen) but fell short of fully correcting this PIT offset.
  • We scanned further the M1 M2 input alignment, and briefly lost the first reflection.
    • We recovered the first reflection, but the beam spot is now incident near the LL OSEM position.
  • We kept scanning M1 - M2 with Ian on the ITMX chamber this time, providing feedback through facetime.
  • Leaning on the ITMX chamber table, we noticed the magnitude of PIT correction left to be done and verified that M1-M2 two axis alignment + ITMX offsets should be enough for it.

We stopped our effort here, the XAUX beam spot is near the lower half of the ITMX face. Tomorrow, we will resume, but we will use airpods and a clean go-pro for real-time audiovisual feedback. Furthermore, ITMX OSEMs should be rebalanced as they haven't been touched after the table was balanced for PR2 and LO1.

ELOG V3.1.3-